Near Earth’s surface, the time t required for an object to fall a distance d is given by t = 1/4d ^1/2 where t is measured in seconds and d is measured in feet. Find the time it will take an object to fall 100 feet.
Answer by stanbon(75887)

Near Earth’s surface, the time t required for an object to fall a distance d is given by t = 1/4d ^1/2 where t is measured in seconds and d is measured in feet. Find the time it will take an object to fall 100 feet.
----------------
t = (1/4)sqrt(d)
-------
t = (1/4)sqrt(100)
---
t = (1/4)10
---
t = 5/2
---
time = 2.5 seconds
--------------------
Cheers,
Stan H.
--------------